OpenStudy (anonymous):

Will give medal and fan for answer. If f(x)=2x^2 +4x -14 and g(x) = -x^3 +15x then what is (f + g)(3)

OpenStudy (anonymous):

-3x^3+6x^2+57x-42

OpenStudy (campbell_st):

well add the 2 functions to find (f + g) so (f + g)(x) = -x^3 +2x^2 +19x -14 now substitute x = 3 to find (f + g)(3) = -(3)^3 + 2(3)^2 + 19(3) - 14 hope it helps